# Scale-Dependent Radial Geometry and Metric Mismatch in Wasserstein Propagation for Reverse Diffusion

### 20 Mar 2026 v1

@@ Abstract
Existing analyses of reverse diffusion often propagate sampling error in the Euclidean geometry underlying $W_{2}$ along the entire reverse trajectory. Under weak log-concavity, however, Gaussian smoothing can create contraction first at large separations while short separations remain non-dissipative. The first usable contraction is therefore radial rather than Euclidean, creating a metric mismatch between the geometry that contracts early and the geometry in which the terminal error is measured. We formalize this mismatch through an explicit radial lower profile for the learned reverse drift. Its far-field limit gives a contraction reserve, its near-field limit gives the Euclidean load governing direct $W_{2}$ propagation, and admissible switch times are characterized by positivity of the reserve on the remaining smoothing window. We exploit this structure with a one-switch routing argument. Before the switch, reflection coupling yields contraction in a concave transport metric adapted to the radial profile. At the switch, we convert once from this metric back to $W_{2}$ under a $p$-moment budget, and then propagate the converted discrepancy over the remaining short window in Euclidean geometry. For discretizations of the learned reverse SDE under $L^{2}$ score-error control, a one-sided Lipschitz condition of score error, and standard well-posedness and coupling hypotheses, we obtain explicit non-asymptotic end-to-end $W_{2}$ guarantees, a scalar switch-selection objective, and a sharp structural limit on the conversion exponent within the affine-tail concave class.
@@ 1 Introduction
Score-based diffusion models are now a central framework for generative modeling, from high-fidelity image synthesis to conditional tasks such as inpainting and super-resolution (Ho et al., 2020; Song et al., 2020; Dhariwal and Nichol, 2021; Rombach et al., 2022). At a conceptual level, sampling follows a reverse-time dynamics. The geometry of the smoothed intermediate laws governs this dynamics, score approximation perturbs it, and numerical samplers discretize it. An end-to-end theory of diffusion sampling should therefore explain, along reverse time, how geometry, initialization mismatch, score error, and discretization interact.
A growing theoretical literature has established convergence guarantees for diffusion samplers in several discrepancy measures, including KL divergence, total variation, and Wasserstein distances (Conforti et al., 2025; Lee et al., 2023; Gao et al., 2025; Gao and Zhu, 2024; Gentiloni-Silveri and Ocello, 2025). We focus on the $2$-Wasserstein distance $W_{2}$, in which initialization mismatch, the perturbation induced by score approximation, and discretization admit a common coupling-based treatment. In globally regular regimes, discrepancy can often be propagated directly in $W_{2}$ over the entire reverse trajectory, without changing metric across phases (Gao et al., 2025; Gao and Zhu, 2024). Our question is: once global regularity is lost, what effective transport geometry becomes operative along the reverse trajectory, how does it depend on time and scale, and how should end-to-end discrepancy be propagated through it?
We work under weak log-concavity: the target law remains globally confining but may be locally nonconvex, in the spirit of recent Ornstein–Uhlenbeck regularization analyses beyond global log-concavity (Gentiloni-Silveri and Ocello, 2025). In this regime, Gaussian smoothing need not yield Euclidean contractivity uniformly across scales. Instead, the lower radial profile we derive can become contractive first at large separations while remaining obstructed near the origin. A standard full-horizon Euclidean propagation is therefore controlled by the short-scale obstruction, even though useful contraction is already present at larger scales. In other words, once global regularity is lost, the effective transport geometry is no longer uniformly Euclidean: it becomes scale-dependent, and this is the metric mismatch behind our analysis.
This leads to a phase-dependent routing argument. On an early window, discrepancy is propagated in a concave transport metric adapted to the radial profile. At a switch point $s_{0}$ in the forward smoothing variable—formalized later through an admissibility condition—we convert once back to $W_{2}$, and the remaining late window is propagated in Euclidean geometry. Section 4 states the resulting one-switch routing theorem for learned reverse diffusion.
@@
    We identify the scale-dependent transport geometry created by Gaussian smoothing under weak log-concavity. An explicit radial lower profile separates a far-field contractive regime from a near-field obstructed regime, and this one-dimensional profile characterizes admissible switch points.
@@
    We turn this geometry into a proof mechanism. On the early window, reflection coupling reduces the separation process of two copies of the learned reverse SDE to a one-dimensional radial diffusion, from which we construct an adapted concave switch metric $W_{\varphi_{s_{0}}}$ (see Definition 6) that propagates initialization mismatch, score forcing, and early discretization up to the switch. We then prove a one-time conversion estimate from $W_{\varphi_{s_{0}}}$ back to $W_{2}$, with exponent $\theta_{p}=(p-2)/(2(p-1))$ under $p$-moment switch control, and show that this exponent is structurally sharp within the affine-tail concave class used for the switch metric.
@@
    We compose these ingredients into an end-to-end one-switch $W_{2}$ routing theorem together with a scalar switch-selection objective, thereby formalizing the separation from the standard direct full-horizon Euclidean route.
@@ 1.1 Technique overview
The proof is driven by a single geometric mechanism. Beyond globally regular regimes, reverse diffusion is not controlled by one uniform Euclidean dissipativity constant, but by a scale-dependent radial contractivity profile of the learned reverse drift. Under weak log-concavity, Gaussian smoothing can make this profile contractive first at large separations while the Euclidean one-sided monotonicity bound relevant for direct $W_{2}$ propagation remains non-contractive at short scales. The key issue is therefore a metric mismatch: the first usable contraction appears in a transport geometry different from the Euclidean geometry in which the terminal error is propagated.
@@
Under Assumptions 1 and 3—the latter being a one-sided geometric control of the score error, rather than a global Lipschitz assumption—the radial profile dictionary in Section 3 yields
@@
The key point is that this single lower envelope already encodes the two geometric regimes used later:
@@
A direct $W_{2}$ propagation uses synchronous coupling. Because the two copies are driven by the same Brownian motion, the noise cancels from the distance evolution, so a Euclidean Grönwall closure requires a global one-sided bound
@@
In globally regular regimes, such a closure is typically provided by a global one-sided Lipschitz control of the learned score or reverse drift (Gao et al., 2025; Gao and Zhu, 2024). The key restriction is that (3) must hold uniformly over all radii. Hence synchronous coupling is governed by the worst value of the radial profile $r\mapsto\kappa_{\widehat{b}_{t}}(r)$, which in the early weakly log-concave regime comes from the near field. So the direct Euclidean route fails not because large-scale contraction is absent, but because synchronous coupling can use it only when every radius is already Euclideanly dissipative.
To exploit the full scale-dependent profile, we instead use reflection coupling (Eberle, 2016; Eberle et al., 2019). Fix an admissible switch $s_{0}$, and compress the early geometry into a switch-level lower envelope $\underline{\kappa}_{s_{0}}$, negative near the origin and positive beyond a threshold $R_{\mathrm{sw}}(s_{0})$. If $Y_{t}$ and $\bar{Y}_{t}$ are reflected copies of the learned reverse SDE and $r_{t}:=\|Y_{t}-\bar{Y}_{t}\|$, then, up to coupling time, the multidimensional dynamics reduce schematically to
@@
Unlike synchronous coupling, the radial dynamics retains a Brownian term and therefore sees the full switch-level profile rather than only its worst short-scale Euclidean value. This is exactly what makes a concave radial cost effective: applying Itô’s formula to $\varphi(r_{t})$ produces the second-order term $2g(T-t)^{2}\varphi^{\prime\prime}(r_{t})$, which can offset the near-field load.
@@
The switch metric $\varphi_{s_{0}}$ is therefore not an auxiliary softer cost chosen after the fact. It is built for the one-dimensional generator exposed by reflection coupling. Writing $\underline{g}(s_{0}):=\inf_{u\in[s_{0},T]}g(u),$ the relevant radial operator has the schematic form
@@
We choose $\varphi_{s_{0}}$ as an explicit concave subsolution of
@@
for switch-dependent parameters $\lambda(s_{0})$ and $\mathfrak{a}(s_{0})$. Its form matches the two ends of the profile. On $0\leq r\leq R_{\mathrm{sw}}(s_{0})$, where $\underline{\kappa}_{s_{0}}$ may still be negative, the curvature term $2\underline{g}(s_{0})^{2}\varphi_{s_{0}}^{\prime\prime}$ supplied by the reflected radial noise compensates the near-field Euclidean obstruction. On $r\geq R_{\mathrm{sw}}(s_{0})$, where the profile carries positive radial margin, no curvature is needed; the metric is continued affinely so as to retain sensitivity to large separations and hence preserve the far-field reserve.
@@
where $P_{u,v}$ denotes the learned reverse semigroup. Thus the far-field reserve is converted into actual early-window transport contraction, but in an adapted concave metric rather than directly in $W_{2}$.
This contraction compresses the entire early window into one damped switch-time quantity through the same one-step defect interface later used in the Euclidean phase. Because $\varphi_{s_{0}}(r)\leq r$, every one-step $W_{2}$ defect also controls the corresponding $W_{\varphi_{s_{0}}}$ defect.
@@
The switch should not be read as the time at which the geometry suddenly becomes Euclidean. Its role is different: it is the point at which an already damped discrepancy is returned once to the Euclidean geometry in which the final error is measured. This return is possible because the switch metric has an affine tail,
@@
so $W_{\varphi_{s_{0}}}$ controls a truncated linear distance. Combined with a $p$-moment budget at switch time, this yields the interpolation
$W_{2}^{2}\;\lesssim\;\frac{\rho^{2}}{\varphi_{s_{0}}(\rho)}\,\Delta_{\varphi}^{\mathrm{sw}}(s_{0})+\frac{\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})}{\rho^{p-2}},$
@@
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),\operatorname{Law}(\widetilde{X}_{t_{s}})\bigr)\;\lesssim\;C_{p}^{\mathrm{sw}}(s_{0})\,\bigl(\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\bigr)^{\theta_{p}},\qquad\theta_{p}=\frac{p-2}{2(p-1)}.$
It is the quantitative price of recovering quadratic Euclidean transport from an early metric that is deliberately softer than $W_{2}$.
The conversion is also where the central tradeoff becomes visible. A more concave early metric is better at absorbing the near-field obstruction under the reflected radial dynamics, but within the affine-tail concave class this comes with weaker quadratic sensitivity at large distances and hence a more expensive return to $W_{2}$. The sharpness result shows that this loss is structural rather than a proof artifact: the same non-Euclidean softness that makes early damping possible also prevents a free recovery of Euclidean quadratic transport.
@@
After the switch, what remains is to propagate the converted discrepancy over the final short window in the target metric $W_{2}$, for which synchronous coupling is again the natural tool. This yields
@@
Admissibility and optimization therefore play different roles. Admissibility is a geometric existence statement: it guarantees that a genuinely non-Euclidean contractive regime is available on the early window. Optimal switching is a routing statement: it decides when the already compressed discrepancy should be re-exposed to $W_{2}$, balancing how much early damping remains available, how expensive the one-time conversion is, and how much late Euclidean load is left.
@@ 2 Related Work
Our work lies at the intersection of Wasserstein convergence theory for score-based generative models, analyses of reverse-time geometry beyond global log-concavity, and reflection-coupling contraction methods for nonconvex diffusions.
@@
A broad theoretical literature studies diffusion sampling under minimal assumptions or in KL/path-space metrics (Chen et al., 2022, 2023a; Lee et al., 2023). Within $W_{2}$, globally regular regimes admit direct full-horizon Euclidean propagation for reverse SDEs and PFODEs (Gao et al., 2025; Gao and Zhu, 2024). More recent $W_{2}$ analyses move beyond global log-concavity, including Gaussian-tail, semiconvex, and weakly log-concave settings (Wang and Wang, 2024; Bruno and Sabanis, 2025; Gentiloni-Silveri and Ocello, 2025). The three closest comparisons emphasize different aspects of time-varying geometry. Gentiloni-Silveri and Ocello (2025) shows that weak log-concavity can produce alternating contractive and non-contractive phases along reverse time; Bruno and Sabanis (2025) studies semiconvex settings with time-dependent Euclidean $W_{2}$ control; and Wang and Wang (2024) obtains Euclidean/PDE-based bounds under Gaussian-tail assumptions. Our result is complementary to these works. The additional point we emphasize is not only when contractivity appears, but also in which metric it first becomes exploitable: a reverse trajectory may already contract in a softer radial transport geometry even while direct Euclidean closure is still blocked by the near field. This is the metric-mismatch principle behind the one-switch route.
@@
Methodologically, we build on the reflection-coupling line for diffusions with nonconvex geometry, from the classical multidimensional diffusion constructions of Lindvall and Rogers (1986); Chen and Li (1989) to the concave-cost contraction framework of Eberle (2016); Eberle et al. (2019). We use the Gaussian-smoothed radial profile of the learned reverse drift to determine an early-window concave transport metric, propagate score and discretization errors on that window, and then perform a single return to $W_{2}$. Our distinctive contribution is therefore the reverse-diffusion interface and one-switch routing theorem, not reflection coupling in isolation.
@@
A complementary line develops convergence analyses for deterministic, stochastic, and higher-order samplers (Chen et al., 2023b; Li et al., 2024; Beyler and Bach, 2025; Yu and Yu, 2025; Pfarr et al., 2026), as well as sharper complexity or stability/forgetting guarantees (Benton et al., 2023; Chen et al., 2026; Strasman et al., 2026). Here the relation is especially direct. Results such as Beyler and Bach (2025) and Pfarr et al. (2026) provide local or global $W_{2}$ discretization estimates for deterministic, stochastic, and higher-order schemes. In our theorem, those estimates populate the one-step defect interface $\xi_{k}$: once $\xi_{k}$ is available, the routing bound determines how much of that sampler error is damped on the early radial window and how much is exposed to the late Euclidean window. Thus better orders or smaller constants from EM, Heun, PF-ODE, or higher-order analyses can be imported directly without changing the geometric part of the proof. Accordingly, our result is not broader sampler coverage or sharper complexity guarantees, but a propagation principle that separates the metric in which contraction first appears from the metric in which the final error is measured.

### 30 Mar 2026 v2

# Wasserstein Propagation for Reverse Diffusion under Weak Log-Concavity: Exploiting Metric Mismatch via One-Switch Routing

@@ Abstract
Existing analyses of reverse diffusion typically propagate sampling error in the Euclidean geometry underlying $W_{2}$ throughout the reverse trajectory. Under weak log-concavity, this can be suboptimal: Gaussian smoothing may create contraction first at large separations, while short-scale Euclidean dissipativity is still absent. We show that exploiting this metric mismatch can yield strictly sharper end-to-end $W_{2}$ bounds than direct full-horizon Euclidean propagation on mismatch windows. Our analysis derives an explicit radial lower profile for the learned reverse drift, whose far-field and near-field limits quantify the contraction reserve and the residual Euclidean load, respectively. This profile determines admissible switch times and leads to a one-switch routing theorem: reflection coupling damps initialization mismatch, pre-switch score forcing, and pre-switch discretization in an adapted concave transport metric; a single $p$-moment interpolation converts the damped switch-time discrepancy back to $W_{2}$; and synchronous coupling propagates the remaining error over the late Euclidean window. Under $L^{2}$ score-error control, a one-sided monotonicity condition on the score error, and standard well-posedness and coupling assumptions, we obtain explicit non-asymptotic end-to-end $W_{2}$ guarantees, a scalar switch-selection objective, and a conversion exponent $\theta_{p}=(p-2)/(2(p-1))$ that cannot be improved uniformly within the affine-tail concave class under the same $p$-moment switch assumption. For a fixed switch, the routed and direct Euclidean bounds share the same late-window term, so any strict improvement is entirely an early-window effect.
@@ 1 Introduction
Score-based diffusion models are now a central framework for generative modeling, building on early diffusion and score-based formulations (Sohl-Dickstein et al., 2015; Song and Ermon, 2019; Ho et al., 2020; Song et al., 2020). At a conceptual level, sampling is a reverse-time transport problem. The smoothed intermediate laws determine its geometry, score approximation perturbs the drift, and numerical samplers discretize the dynamics. An end-to-end theory of diffusion sampling should therefore identify the geometry that actually governs error propagation along reverse time and turn it into sharp non-asymptotic bounds.
A growing theoretical literature provides convergence guarantees for diffusion samplers in several discrepancy measures, including KL divergence, total variation, and Wasserstein distances (Conforti et al., 2025; Lee et al., 2023; Gao et al., 2025; Gao and Zhu, 2024; Gentiloni-Silveri and Ocello, 2025). We focus on the $2$-Wasserstein distance $W_{2}$, in which initialization mismatch, score approximation, and discretization admit a common coupling-based treatment. In globally regular regimes, discrepancy can often be propagated directly in $W_{2}$ along the entire reverse trajectory, without changing metric across phases (Gao et al., 2025; Gao and Zhu, 2024). Once global regularity is lost, however, this Euclidean route need not remain sharp. The natural question is: what transport geometry actually governs the propagation of sampling error along the reverse trajectory, and how should end-to-end $W_{2}$ bounds reflect that geometry rather than a uniform Euclidean worst case?
We study this question under weak log-concavity, where the target law remains globally confining but may be locally nonconvex, in the spirit of recent Ornstein–Uhlenbeck regularization analyses beyond global log-concavity (Gentiloni-Silveri and Ocello, 2025). In this regime, Gaussian smoothing need not produce Euclidean contractivity uniformly across scales. Instead, the radial lower profile we derive may become contractive first at large separations while remaining obstructed near the origin. A direct full-horizon Euclidean propagation is then bottlenecked by the short-scale obstruction even though contraction is already available farther out. The effective transport geometry is therefore scale-dependent, and treating it as uniformly Euclidean can make the resulting end-to-end $W_{2}$ bound genuinely conservative.
On an early window, discrepancy is propagated in a concave transport metric adapted to the radial profile. At an admissible switch point $s_{0}$ in the forward smoothing variable, the damped discrepancy is returned once to $W_{2}$, and the remaining late window is propagated in Euclidean geometry. For a fixed admissible switch, the routed and direct Euclidean bounds share the same late-window term, so any advantage is entirely an early-window effect; on mismatch windows, the routed bound can be strictly sharper.
@@
    We derive an explicit radial lower profile for the learned reverse drift under weak log-concavity. This profile separates far-field contraction from near-field obstruction and yields a transparent admissibility criterion for switch points.
@@
    We develop the early-window contraction mechanism behind the routed analysis. Reflection coupling reduces the separation process of two copies of the learned reverse SDE to a one-dimensional radial diffusion, from which we construct an adapted concave switch metric $W_{\varphi_{s_{0}}}$ that propagates initialization mismatch, pre-switch score forcing, and early discretization up to the switch.
@@
    We prove an end-to-end one-switch $W_{2}$ routing theorem. The theorem localizes metric mismatch at a single interface, returns the pre-switch discrepancy from $W_{\varphi_{s_{0}}}$ to $W_{2}$ only once under a $p$-moment budget, induces a scalar switch-selection objective over admissible grid-aligned switches, and yields the exponent $\theta_{p}=(p-2)/(2(p-1))$, which cannot be improved uniformly within the affine-tail concave class under the same switch-moment assumption.
  4. (iv)
    We give an exact comparison with the standard direct full-horizon Euclidean route. For any fixed admissible switch, the routed and direct bounds share the same late-window term, so the comparison reduces entirely to the early window. This identifies mismatch windows on which the routed bound can be strictly sharper.
@@ 1.1 Technique overview
The proof is organized around a single geometric fact. Beyond globally regular regimes, reverse diffusion is controlled not by one uniform Euclidean dissipativity constant but by a scale-dependent radial contractivity profile of the learned reverse drift. Under weak log-concavity, Gaussian smoothing can make this profile contractive first at large separations even while the Euclidean one-sided monotonicity bound required for direct $W_{2}$ propagation remains non-contractive near the origin. The first usable contraction therefore appears in a transport geometry different from the Euclidean geometry in which the final error is measured.
@@
Under Assumptions 1 and 3—the latter being a one-sided monotonicity condition on the score error rather than a global Lipschitz assumption—the radial profile dictionary in Section 3 yields
@@
A single lower envelope controls both geometric regimes used later:
@@
A direct $W_{2}$ propagation uses synchronous coupling. Because the two copies are driven by the same Brownian motion, the noise cancels in the distance dynamics, so a Euclidean Grönwall closure requires a global one-sided bound
@@
In globally regular regimes, such a closure is typically obtained from a global one-sided Lipschitz control of the learned score or reverse drift (Gao et al., 2025; Gao and Zhu, 2024). The restriction is that (3) must hold uniformly over all radii. Synchronous coupling is therefore governed by the worst part of the radial profile $r\mapsto\kappa_{\widehat{b}_{t}}(r)$, which in the early weakly log-concave regime is precisely the near field. The direct Euclidean route thus fails not because large-scale contraction is absent, but because it cannot exploit that contraction before every radius is already Euclideanly dissipative.
To use the full scale-dependent profile, we instead employ reflection coupling (Eberle, 2016; Eberle et al., 2019). Fix an admissible switch $s_{0}$, and compress the early geometry into a switch-level lower envelope $\underline{\kappa}_{s_{0}}$, negative near the origin and positive beyond a threshold $R_{\mathrm{sw}}(s_{0})$. If $Y_{t}$ and $\bar{Y}_{t}$ are reflected copies of the learned reverse SDE and $r_{t}:=\|Y_{t}-\bar{Y}_{t}\|$, then, up to coupling time, the multidimensional dynamics reduce schematically to
@@
Unlike synchronous coupling, the radial dynamics retains a Brownian term and therefore sees the entire switch-level profile rather than only its worst short-scale Euclidean value. This is exactly why a concave radial cost becomes effective: applying Itô’s formula to $\varphi(r_{t})$ produces the second-order term $2g(T-t)^{2}\varphi^{\prime\prime}(r_{t})$, which can offset the near-field load.
@@
The switch metric $\varphi_{s_{0}}$ is therefore not an auxiliary softer cost chosen after the fact. It is the radial Lyapunov gauge naturally associated with the reflected generator. Writing $\underline{g}(s_{0}):=\inf_{u\in[s_{0},T]}g(u),$ the relevant radial operator has the schematic form
@@
We construct $\varphi_{s_{0}}$ as an explicit concave subsolution of
@@
for switch-dependent parameters $\lambda(s_{0})$ and $\mathfrak{a}(s_{0})$. Its form mirrors the two ends of the profile. On $0\leq r\leq R_{\mathrm{sw}}(s_{0})$, where $\underline{\kappa}_{s_{0}}$ may still be negative, the curvature term $2\underline{g}(s_{0})^{2}\varphi_{s_{0}}^{\prime\prime}$ supplied by the reflected radial noise compensates the near-field obstruction. On $r\geq R_{\mathrm{sw}}(s_{0})$, where the profile carries positive radial margin, no curvature is needed; the metric is continued affinely to preserve sensitivity to large separations and hence retain the far-field reserve.
@@
where $P_{u,v}$ denotes the learned reverse semigroup. Thus the far-field reserve is converted into genuine early-window transport contraction, but in an adapted concave metric rather than directly in $W_{2}$.
This contraction compresses the entire early window into a single damped switch-time quantity through the same one-step defect interface later used in the Euclidean phase. Since $\varphi_{s_{0}}(r)\leq r$, one has
$W_{\varphi_{s_{0}}}(\mu,\nu)\leq W_{1}(\mu,\nu)\leq W_{2}(\mu,\nu),$
so each one-step $W_{2}$ defect also controls the corresponding $W_{\varphi_{s_{0}}}$ defect.
@@
The switch should not be read as the time at which the geometry suddenly becomes Euclidean. Its role is different: it is the point at which an already damped discrepancy is returned once to the Euclidean metric in which the final error is measured. This return is possible because the switch metric has a global linear lower bound,
@@
and hence
$W_{1}(\mu,\nu)\leq\mathfrak{a}(s_{0})^{-1}W_{\varphi_{s_{0}}}(\mu,\nu).$
Combined with a $p$-moment budget at switch time, this yields the interpolation
$W_{2}^{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)\;\lesssim\;\frac{\rho}{\mathfrak{a}(s_{0})}\,\Delta_{\varphi}^{\mathrm{sw}}(s_{0})+\frac{\overline{\mathcal{M}}_{p}^{\mathrm{sw}}(s_{0})}{\rho^{p-2}},$
@@
$W_{2}\!\bigl(\operatorname{Law}(Z_{t_{s}}),p_{s_{0}}\bigr)\;\lesssim\;C_{p}^{\mathrm{sw}}(s_{0})\,\bigl(\Delta_{\varphi}^{\mathrm{sw}}(s_{0})\bigr)^{\theta_{p}},\qquad\theta_{p}=\frac{p-2}{2(p-1)}.$
This is the only place where the exponent $\theta_{p}$ enters. It measures the price of recovering quadratic Euclidean transport from a switch metric that is deliberately softer than $W_{2}$.
The conversion is also where the central tradeoff becomes explicit. A more concave early metric is better at absorbing the near-field obstruction under the reflected radial dynamics, but within the affine-tail concave class this comes with weaker quadratic sensitivity at large distances and hence a more expensive return to $W_{2}$. The sharpness result shows that this loss is structural rather than a proof artifact: the same non-Euclidean softness that makes early damping possible also prevents a free recovery of Euclidean quadratic transport.
@@
After the switch, the remaining interval is propagated in the target metric $W_{2}$, for which synchronous coupling is again the natural tool. This yields
@@
Admissibility and optimization therefore play different roles. Admissibility is a geometric existence statement: it certifies that a genuinely non-Euclidean contractive regime is available on the early window. Optimal switching is a routing statement: it decides when the already compressed discrepancy should be re-exposed to $W_{2}$, balancing the remaining early damping, the one-time conversion cost, and the residual late Euclidean load.
@@ 2 Related Work
Our work lies at the intersection of reverse-time diffusion representations (Anderson, 1982; Haussmann and Pardoux, 1986), Wasserstein convergence theory for score-based generative models, analyses of reverse-time geometry beyond global log-concavity, and reflection-coupling contraction methods for nonconvex diffusions.
@@
A broad theoretical literature studies score-based diffusion sampling under minimal assumptions, polynomial-complexity regimes, manifold-supported data, and KL/path-space metrics (Azangulov et al., 2024; Conforti et al., 2025; Chen et al., 2022, 2023a; Lee et al., 2023). Within Wasserstein control, strongly log-concave and globally regular settings admit direct Euclidean propagation arguments for reverse SDEs and PF-ODEs (Bruno et al., 2025; Gao et al., 2025; Gao and Zhu, 2024). More recent $W_{2}$ analyses move beyond global log-concavity, including Gaussian-tail, semiconvex, weak-semiconvex, and weakly log-concave regimes (Wang and Wang, 2024; Conforti, 2024; Bruno and Sabanis, 2025; Gentiloni-Silveri and Ocello, 2025; Kremling et al., 2025). The closest comparisons from our viewpoint emphasize different manifestations of time-varying geometry. Gentiloni-Silveri and Ocello (2025) shows that weak log-concavity can produce alternating contractive and non-contractive phases along reverse time; Bruno and Sabanis (2025) studies semiconvex settings with time-dependent Euclidean $W_{2}$ control; and Kremling et al. (2025) proves weak-log-concave guarantees for PF-ODEs under Euclidean/Lipschitz control. Our result is complementary to these works. The additional point we emphasize is not only when contractivity appears, but also in which metric it first becomes exploitable: a reverse trajectory may already contract in a softer radial transport geometry even while direct Euclidean closure is still blocked by the near field. This metric-mismatch principle is what leads to our one-switch route and is also what allows the routed bound to be sharper than direct full-horizon Euclidean propagation on mismatch windows.
@@
Methodologically, we build on the reflection-coupling program for diffusions with nonconvex geometry, from the classical multidimensional constructions of Lindvall and Rogers (1986); Chen and Li (1989) and the coupling overview of Kallenberg (1993) to the concave-distance contraction framework of Eberle (2016); Eberle et al. (2017, 2019). Because our pre-switch analysis compares learned and exact reverse dynamics with different drifts, the sticky-coupling analysis of Eberle and Zimmer (2016) is also methodologically relevant. We use the Gaussian-smoothed radial profile of the learned reverse drift to determine an early-window concave transport metric, propagate initialization mismatch, score forcing, and discretization errors on that window, and then perform a single return to $W_{2}$. Our distinctive contribution is therefore the reverse-diffusion interface and one-switch routing theorem, not reflection coupling in isolation.
@@
A separate line analyzes deterministic, stochastic, and higher-order diffusion samplers (Chen et al., 2023b; Li et al., 2024; Beyler and Bach, 2025; Yu and Yu, 2025; Pfarr et al., 2026), together with sharper complexity and stability guarantees (Benton et al., 2023; Chen et al., 2026; Strasman et al., 2026). Our theorem is modular with respect to those discretization results. Any one-step or accumulated $W_{2}$ defect estimate enters only through the interface term $\xi_{k}$, after which the routing theorem determines how much of that sampler error is damped on the early radial window and how much is exposed to the late Euclidean window. Thus our contribution is not a new sampler analysis per se, but a propagation theorem that can incorporate sharper sampler bounds without changing the geometric argument.